# Responsible AI Question Bank: A Comprehensive Tool for AI Risk Assessment

### v1

@@ 6. Does the model provider provide intelligible instructions for use? User instructions
Effective data governance necessitates that model providers rigorously evaluate the quality of training data sources to ensure they are suitable and free from bias. It is also critical to maintain thorough documentation of any training data protected by copyright law. Key questions to consider can include Does the model provider assess the quality of the data sources used for training? and Does the model provider document detailed summaries of the use of training data protected under copyright law? Evidence of adherence to these standards is found in detailed summaries of the training data employed. As known solutions to support this, Datasheets for datasets [29] and Data bill of materials can be considered.
@@
The foundation model providers should assess and document the sustainability of AI model training and management activities. This includes measuring factors such as energy usage, carbon emissions, and tons of waste generated. Model providers should also consider additional metrics related to model/data size, complexity, training time, and computational performance (e.g., FLOPs), as these factors can indirectly impact the environment. Providing detailed assessments and documentation of these metrics ensures transparency and accountability in their environmental practices.
Quality management is essential for model providers. It mandates rigorous evaluations and meticulous documentation throughout the lifecycle of AI models. This process entails thorough testing and analysis to ensure reliability, accuracy, and adherence to predefined performance criteria. The testing report serves as crucial evidence, detailing the evaluation methods used, findings, and any adjustments made over the lifecycle. These practices enable the model providers to enhance transparency and facilitate continuous improvement in their model development processes, thereby allowing agent developers to ascertain the model’s trustworthiness.
The main purpose of this section is to demonstrate how our RAI Question Bank can be used for the risk assessment of AI agent systems, rather than developing a comprehensive risk assessment framework. This section, therefore, may lack detailed methodologies or extensive examples, as its focus is on illustrating the practical application of the question bank.
@@ 7 Conclusion
The RAI Question Bank represents a significant step forward in the structured risk assessment of AI systems. Its hierarchical structure, organized into different levels (Level 1-3), ensures a comprehensive evaluation by linking high-level principles with detailed, technical questions. By organizing risk themes and questions around eight AI ethics principles, it helps effectively identify potential risks, ensure compliance with emerging regulations, and enhance overall AI governance.
@@
In conclusion, the RAI Question Bank offers a valuable resource and tool for organizations by promoting a thorough and structured approach to risk assessment. It supports the responsible use and development of AI systems that are not only innovative but also aligned with ethical principles and societal values.

### v2

@@ 6. Does the model provider provide intelligible instructions for use? User instructions
Effective data governance necessitates that model providers rigorously evaluate the quality of training data sources to ensure they are suitable and free from bias. It is also critical to maintain thorough documentation of any training data protected by copyright law. Evidence of adherence to these standards is found in detailed summaries of the training data employed. As known solutions to support this, Datasheets for datasets [106] and Data bill of materials can be considered.
@@
The foundation model providers should assess and document the sustainability of AI model training and management activities. This includes measuring factors such as energy usage, carbon emissions, and tons of waste generated [107]. Model providers should also consider additional metrics related to model/data size, complexity, training time, and computational performance (e.g., FLOPs), as these factors can indirectly impact the environment. Providing detailed assessments and documentation of these metrics ensures transparency and accountability in their environmental practices.
Quality management mandates rigorous evaluations and meticulous documentation throughout the lifecycle of AI models [21]. This process entails thorough testing and analysis to ensure reliability, accuracy, and adherence to predefined performance criteria. The testing report serves as crucial evidence, detailing the evaluation methods used, findings, and any adjustments made over the lifecycle. These practices enable the model providers to enhance transparency and facilitate continuous improvement in their model development processes, thereby allowing agent developers to ascertain the model’s trustworthiness.
@@ 7 Conclusion
In this study, we presented the RAI Question Bank, developed through comprehensive approaches, including a systematic literature review, an in-depth survey of existing RAI frameworks, and multiple case studies. The RAI Question Bank represents a significant step forward in the structured risk assessment of AI systems. Its hierarchical structure, organized into different levels (Level 1-3), ensures a comprehensive evaluation by linking high-level principles with detailed, technical questions. By organizing RAI themes and risk questions around eight AI ethics principles, it helps effectively identify potential risks, ensure compliance with emerging regulations, and enhance overall AI governance.
@@
In conclusion, the RAI Question Bank offers valuable insights and serves as a knowledge base, functioning as a core resource and tool for organizations, while promoting a thorough and structured approach to risk assessment. It supports the responsible use and development of AI systems that are not only innovative but also aligned with ethical principles and societal values.